OpenStudy (anonymous):

How can I resolve this equation by substitution? \[\int\limits{(x^2)/1-x^3}dx\]

OpenStudy (bahrom7893):

let u = (1-x^3)

OpenStudy (experimentx):

let x^3 = y, then 3x^2 = dy

OpenStudy (bahrom7893):

No let u = 1-x^3

OpenStudy (bahrom7893):

du=-3x^2dx

OpenStudy (bahrom7893):

Multiply your integral by: (-1/3)*(-3)

OpenStudy (experimentx):

1/3*int 1/(1-y)*dy

OpenStudy (bahrom7893):

So you'll end up with: (-1/3)Integral((1/u)du)

OpenStudy (bahrom7893):

my method is easier experimentX

OpenStudy (experimentx):

yeah ... faster too

OpenStudy (anonymous):

both is nice, thank you...

OpenStudy (bahrom7893):

(-1/3)Ln|u|+C=(-1/3)Ln|1-x^3|+C

OpenStudy (anonymous):

Thanks